OpenStudy (anonymous):

how do i simplify 16/28

OpenStudy (anonymous):

find the largest common denominator

OpenStudy (anonymous):

To simplify or reduce 16/28, first recognize that their greatest common factor is 4. Divide both by 4 for your answer: 4/7.

OpenStudy (austinl):

If you want to be really simplistic, you can just see that they are even numbers, and just divide by 2 until you cannot divide anymore.

OpenStudy (anonymous):

I dont know how to divide

OpenStudy (austinl):

\[16 \div 2 =?\] How many times does 2 go into 16?

OpenStudy (anonymous):

8 times

OpenStudy (austinl):

Exactly. \[16\div 2 = 8\div 2 = 4\] \[28\div 2 = 14\div 2 = 7\] \[\frac{4}{7}\] You have to keep the number of divisions equal too.

OpenStudy (anonymous):

Ok thanks i am a little confused on how to do 6/54

OpenStudy (austinl):

\[54\div 6 =?\]

OpenStudy (anonymous):

9

OpenStudy (austinl):

Correct. 6/6=1 So we have. \[\frac{\frac{6}{6}}{\frac{54}{6}}=?\]

OpenStudy (anonymous):

Am confused

OpenStudy (austinl):

\[\frac{6\div 6}{54\div 6}=?\]

OpenStudy (anonymous):

1/

OpenStudy (anonymous):

1/9

OpenStudy (austinl):

Correct!

OpenStudy (anonymous):

Ok so now how do i do 18/48

OpenStudy (austinl):

\[\frac{18\div6}{48\div6}\]

OpenStudy (anonymous):

Were did you get 6 from

OpenStudy (austinl):

Greatest common factor.

OpenStudy (anonymous):

Ok is 3/8

OpenStudy (anonymous):

Ok i need help with 15/39 but i think is 1/2 am not sure

OpenStudy (austinl):

\[\frac{15\div 3}{39\div 3}\]

OpenStudy (anonymous):

5/13

OpenStudy (anonymous):

How about 16/46